OpenStudy (anonymous):

What is 8x-2y=48 when y =4

OpenStudy (anonymous):

7

OpenStudy (anonymous):

So if y = 4 in this equation, first, plug it in. So you now have 8x - 2(4) = 48. So 2 * 4 is 8 which means you now have 8x - 8 = 48. 48 + 8 = 56 so it's 7.